find the slope of the line containing (3,5) and (1,8)

Answer by sophxmai(62)

Your points are (3,5) and (1,8)
Use the slope formula: m=(y2-y1)/(x2-x1)

m=(8-5)/(1-3)
m=3/-2
m=-3/2